Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

In a room filled with 7 people, 4 people have exactly 1

Expert replies
by varun289 » Fri Dec 14, 2012 12:51 am
233. In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
Join the discussion
Source: — Problem Solving |

by puneetkhurana2000 » Fri Dec 14, 2012 1:06 am
Take the case as 1-2(one sibling each) and 3-4(one sibling each) and 5-6-7(two siblings each) ..so 1-2-3-4 have one sibling each and 5-6-7 have two siblings each.

Total ways 7C2 = 21

Favorable ways 2C1*2C1 + 3C1*2C1 + 3C1*2C1 = 16

Answer is [spoiler]16/21[/spoiler]
Join the discussion

by Brent@GMATPrepNow » Fri Dec 14, 2012 5:51 am
varun289 wrote:233. In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
First we need to recognize that the given information tells us that the 7 people consist of:
- a sibling trio
- a sibling pair
- and another sibling pair

Using counting techniques:

For this question, it's easier to find the complement.
So P(not siblings) = 1 - P(they are siblings)

P(they are siblings) = [# of ways to select 2 siblings] / [total # of ways to select 2 people]

# of ways to select 2 siblings
Case a) 2 siblings from the sibling trio: from these 3 siblings, we can select 2 siblings in 3C2 ways (3 ways)
Case b) 2 siblings from first sibling pair: we can select 2 siblings in 2C2 ways (1 way)
Case c) 2 siblings from second sibling pair: we can select 2 siblings in 2C2 ways (1 way)

So, total number of ways to select 2 siblings = 3+1+1 = 5

total # of ways to select 2 people
We have 7 people and we want to select 2 of them
We can accomplish this in 7C2 ways (21 ways)

So, P(they are siblings) = 5/21

This means P(not siblings) = 1 - 5/21
= [spoiler]16/21[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by ritind » Wed Dec 19, 2012 1:05 am
@Varun : whats the source of your quant questions. Help will be highly appreciated.
Join the discussion

by ritind » Tue Dec 25, 2012 1:55 am
puneetkhurana2000 wrote:Take the case as 1-2(one sibling each) and 3-4(one sibling each) and 5-6-7(two siblings each) ..so 1-2-3-4 have one sibling each and 5-6-7 have two siblings each.

Total ways 7C2 = 21

Favorable ways 2C1*2C1 + 3C1*2C1 + 3C1*2C1 = 16
Answer is [spoiler]16/21[/spoiler]
Can u explain the highlighted part
Join the discussion

by GMATGuruNY » Tue Dec 25, 2012 4:40 am
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by puneetkhurana2000 » Tue Dec 25, 2012 6:19 am
puneetkhurana2000 wrote:
Take the case as 1-2(one sibling each) and 3-4(one sibling each) and 5-6-7(two siblings each) ..so 1-2-3-4 have one sibling each and 5-6-7 have two siblings each.

Total ways 7C2 = 21

Favorable ways 2C1*2C1 + 3C1*2C1 + 3C1*2C1 = 16
Answer is 16/21

ritind wrote:
Can u explain the highlighted part
2C1*2C1(selecting 1 from (1,2) and selecting 1 from (3,4).
3C1*2C1(selecting 1 from (5,6,7) and selecting 1 from (3,4).
3C1*2C1(selecting 1 from (5,6,7) and selecting 1 from (1,2).

Hope this helps!!!
Join the discussion

by ritind » Tue Dec 25, 2012 11:55 pm
GMATGuruNY wrote:I posted an explanation here:

https://www.beatthegmat.com/siblings-t66114.html
Thanks a lot Mitch...this is exactly what I was looking for
Join the discussion